# From Challenges and Pitfalls to Recommendations and Opportunities: Implementing Federated Learning in Healthcare

### v1

@@ 1 Introduction
The rest of this review is organized as follows. Section 2 offers the background and preliminaries of FL. Section 3 presents the screening procedure of this review. Section 4 summarizes the key results of this review. Section 5 discusses the most recent advances, challenges and pitfalls for implementing FL in healthcare, providing recommended practices to avoid current pitfalls as well as outlining potential future research directions and opportunities.
@@ 2 Preliminaries
Beyond centralized or decentralized topologies, FL has evolved to address complex scenarios caused by various feature and sample distributions, leading to the development of Horizontal Federated Learning (HFL), where the data from different clients has significant overlap in the feature space but little overlap in the sample space; Vertical Federated Learning (VFL), where the data from different clients has very little overlap in the feature space while significant overlap in the sample space; and Federated Transfer Learning (FTL), which leverages knowledge transfer to manage scenarios where clients have little overlap in both feature and sample spaces. Figure 2 provides a clear illustration of the differences among HFL, VFL, and FTL. HFL is the most common scenario in FL studies. For example, a vast number of people use wearables, such as the Apple Watch, to monitor their health conditions. These sensors generate massive amounts of data that share the same feature space, which can be used for collaborative model training. See Clients 1 and 2 in Figure 2 for an example. VFL aligns well with real-world applications. For instance, a patient may have medical records at different hospitals, each containing distinct features. With many samples, these hospitals will possess more comprehensive feature sets, enabling them to jointly learn the model more accurately. Refer to Figure 2 to see Clients 1 and 3 for an example. FTL is also well-suited for healthcare scenarios, especially when dealing with tabular EHR data 28. For example, see Clients 2 and 4 in Figure 2, where there is only a small intersection in the feature space, different clients have varying combinations of healthcare data and patients.
FL can be further categorized into cross-silo FL and cross-device FL based on the scale and attributes of participating clients. Cross-silo FL is tailored for scenarios where a limited number of participating clients, such as various hospitals, medical centres, and institutions, engage in all stages of FL training. Cross-device FL, on the other hand, is designed for scenarios with a multitude of participating clients, typically devices with limited data storage capabilities, such as wearables and Internet of Medical Things (IoMT) devices.
@@ 3 Method
Our review was conducted in accordance with the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines.As shown in Figure 3, the flow diagram outlines the search, inclusion and exclusion procedures. We carried out a comprehensive search of the most recent studies focusing on advanced FL technologies within healthcare domain up to May 2024.
@@ 4.1 Application and Data
Included studies explored a broad range of healthcare specialties, including general medicine 6, cardiology 29, oncology 30, ophthalmology 31, drug discovery 15, multiomics 32, dermatology 33, and radiology 34. Most studies focused on tasks such as classification (67/107), segmentation (20/107), and detection (8/107), with additional applications in regression 35, clustering 32, reconstruction 36, 37, feature selection 38, 39, data synthesis 40, 41, 42, and biomedical language process 43. In terms of data types, medical imaging, including MRI, CT, and X-rays, was the most frequently used (55/107), followed by clinical data and electronic health records (EHR) (24/107), skin images (6/107), retinal images (6/107), histopathology slides (16/107), multiomics data (3/107), and biomedical language data (1/107). Some studies involved multiple data types, while 8 studies did not specify the type of data used or used non-healthcare data 44, 45, 46.
@@
Cohort Size $\leq 100$ $100-1000$ $\geq 1000$ unavailable 6 9 24 68 23, 47, 48 5, 6, 7 8, 35, 14 49, 50, 36
Task Classification Segmentation Detection Others 74 22 8 10 51, 52, 53 54, 55, 56 19, 57, 58 35, 32, 37
Data Type Medical Imaging (MRI, CT, X-rays) Clinical and EHR Skin Images Retinal Images Histopathology Slides Multiomics Biomedical Language unavailable or non-healthcare 55 24 6 6 16 3 1 8 59, 6, 8 60, 61, 62 49, 63, 64 65, 63, 56 24, 66, 67 32, 68, 69 43 70, 46, 45
Number of Clients $\leq 10$ $10-50$ $\geq 50$ unavailable 57 17 6 27 7, 11, 15 5, 8, 9 61, 70, 71 72, 73, 46
Topology Centralized Decentralized 95 12 51, 52, 36 74, 75, 44
Type of Federation Cross-Silo Cross-Device unavailable 98 7 2 76, 37, 56 75, 74, 77 46, 73
Scenario Deployment Simulation 10 97 78, 79, 80 81, 82, 64
Framework Custom-designed Open-source Options unavailable 78 13 16 82, 81, 83 13, 8, 79 50, 64, 84
Data Curation Standardization & Harmonization unavailable 12 95 7, 66, 23 49, 85, 86
Data Partition HFL VFL FTL unavailable 94 3 1 12 87, 88, 89 28, 48, 36 77 40, 90, 39
Natural Split Simulate unavailable 37 67 3 75, 91, 37 92, 43, 86 93, 79, 9
Train/Test/Val Details per Client 17 43, 89, 94
Holdout Cohort for Evaluation 12 49, 37, 66
Model Deep Learning Traditional Machine Learning unavailable 83 7 17 78, 80, 57 7, 35, 62 95, 13, 79
Initialization: Random Initialization: Pretrained/Foundation Models unavailable 18 5 84 7, 28, 65 96, 33, 43 97, 83, 53
@@
Generalization in Open Domain 15 7, 28, 33
Communication Efficiency 19 61, 27, 98
@@
Temporal Data Dynamics 2 57, 28
Synchronous Aggregation Asynchronous Aggregation 92 15 19, 23, 24 32, 33, 34
Privacy and Security Model Updates Protection 41 19, 99, 64
Open Source Code Available Trained Model Available unavailable 29 1 78 51, 52, 65 8 100, 47, 98
@@ 4.2 Topology, Scenario and Framework
The centralized FL paradigm dominates current implementations, with 95 out of 107 studies following this topology. Only 10 studies reported real-world deployments in distributed clinical settings, while the rest remained in the realm of prototypes or simulations. In terms of frameworks, the majority (78/107) utilized custom-designed FL frameworks, while a smaller number (13/107) employed open-source options such as Flower 13, Flare 8, 79, 11, SubstraFL 7, 15, TFF 35, OpenFL 14, PySyft 34, 47, 101, 66, and FedBioMed 80. 16 studies did not specify the framework used. Further details about open-source frameworks can be found in Table 2.
@@ 4.3 Data Curation and Partition
Among the reviewed studies, only 12 provided details on the processes of data standardization and harmonization. Regarding data partition, HFL was the predominant approach, with 94 out of 107 studies focusing solely on it. In contrast, VFL was explored in only 3 studies 36, 28, 48, while 2 studies considered both HFL and VFL in combination 28, 48. Only 1 study discussed FTL 77. Notably, 12 studies did not mention this aspect at all. The majority of studies addressed only one type of data heterogeneity, such as quantity skew or label skew, without considering multiple factors simultaneously. Moreover, 37 studies employed natural data splits for training and/or evaluation, while the rest relied on artificial splits. Only 17 studies detailed their training, testing or validation sets and 12 studies split a holdout cohort for evaluation.
@@ 4.4 Model
Among reviewed studies, Convolutional Neural Networks (CNNs) were the most commonly utilized (80/107), including both custom models specifically designed for healthcare tasks and well-established architectures like ResNet, DenseNet, MobileNet, and U-Net 100, 47, 98. Additionally, Recurrent Neural Networks (RNNs) have been incorporated to leverage their strengths in handling complex healthcare data 48, 57, 102. Some studies also employed custom Multi-Layer Perceptrons (MLPs) and attention mechanisms to further boost model performance 70, 103, 61. 43 utilized large language models for distributed biomedical natural language processing. Beyond deep learning approaches, several studies explored traditional machine learning (ML) algorithms, including linear models and ensemble methods. Notable examples include logistic regression 70, support vector machines 84, fuzzy clustering 104, and decision trees 71, 105. Interestingly, only 23 studies explicitly discussed their initialization strategies for model training. Among these, the majority opted for random initialization, while a mere five clearly stated that they utilized pretrained or foundation models as their starting point 96, 106, 33, 29, 43.
@@ 4.5 Optimization
Most studies addressed either data or model heterogeneity, and none of them considered system heterogeneity. Only 15 studies evaluated model generalization ability in unseen open domains. A total of 19 studies focused on improving communication efficiency, employing techniques such as knowledge distillation 51, 99, 96, gradient quantization 61, one-shot FL 98, split learning 101, and tensor factorization 95, 107. In terms of convergence analysis, a few studies (21/107) reported metrics such as communication rounds and costs, as well as overall convergence time, but none provided a theoretical understanding of convergence dynamics. Only two studies considered temporal data dynamics in model learning 28, 57. Regarding synchronization, 15 studies employed asynchronous aggregation instead of synchronous aggregation, particularly in applications involving wearables 75 and IoMT devices 57.
@@ 4.6 Privacy and Security
Only 41 studies addressed the exchange of model updates with privacy guarantees. The most commonly used techniques for safeguarding model updates included Differential Privacy (DP), Homomorphic Encryption (HE), Secure Multi-Party Computation (SMPC), knowledge distillation, and partial model exchange. However, metadata such as sample sizes and distributions were frequently shared without protection, particularly in methods based on FedAvg 106, 108. To mitigate the risk of adversaries inferring raw data, synthetic data was employed in some cases 19, 98, 42, 85. Additionally, swarm learning and blockchain were utilized to secure the communication process 24, 19.
@@ 4.7 Fairness and Incentive
Only three studies have discussed issues related to fairness and/or incentives in healthcare FL 109, 45, 53, with just one of these studies specifically exploring the complexities of both fairness and incentives in detail 53.
@@ 4.8 Evaluation
Most studies used conventional ML metrics for evaluation, such as accuracy, precision, area under the receiver operating characteristic curve (AUC), sensitivity/recall, specificity, F1-score, Dice score, Intersection over Union (IoU), Hausdorff Distance (HD), and loss value. Additionally, many studies performed comparisons against classical centralized models or localized models, and conducted ablation studies. However, only a few studies (26/107) addressed critical aspects unique to FL, such as communication overhead, resource consumption, scalability, generalization, privacy, fairness, and security concerns. As for benchmarking, just one study provided relatively comprehensive benchmarks across multiple healthcare datasets 52. Interpretability was explored in seven studies, either through feature selection 62, 38, attention maps 39, 51, 66, or tree-based models 7, 35. While 29 studies released their source code, only one also made the trained model publicly available 8.
@@ 5.1.1 Domination of Simulation Studies
Moreover, very few studies have explored FL practices at a national or international scale 8, 62. Notable examples include the Collaborative Data Analysis (CODA) 78, 15, HealthChain 12 and DRAGON 13. CODA tested FL’s feasibility across eight hospitals in Canada by enrolling patients with suspected or confirmed COVID-19 over three years. Melody deployed multi-task FL among 10 pharmaceutical companies to optimize the drug discovery process. HealthChain and DRAGON implemented FL across multiple hospitals in Europe, facilitating the prediction of treatment responses for cancer and COVID-19 patients.
@@
Flare 79 Nvidia DP, HE - -
FedML 110 TensorOpera DP, HE -
FederatedScope 111 Alibaba DP, SMPC - - - - - -
Flower 112 FlowerLab DP - -
FATE 113 WeBank DP, HE, SMPC - -
SubstraFL 114 Owkin DP, SMPC - -
PySyft 115 OpenMined DP - - - - - - -
OpenFL 116 Intel DP - - - - - -
TFF 117 Google DP - - - - - -
Fed-BioMed 80 Inria DP - - - - -
IBM FL 118 IBM DP, SMPC - - - - - -
PaddleFL 119 Baidu DP, SMPC - - - - - -
SAFEFL 120 ENCRYPTO SMPC - - - - - - -
@@ 5.1.2 Deficiency in FL Frameworks Development and Usage
Researchers aiming to swiftly prototype and test novel concepts can benefit from frameworks that incorporate comprehensive simulator modules. These tools allow for the rapid iteration and validation of ideas within a controlled simulation environment, which can be critical for the initial stages of research and development. Engineers seeking to deploy FL in real-world scenarios should consider frameworks tailored to the specific needs of the healthcare domain. These frameworks should offer healthcare specific adaptations to ensure compatibility with medical data formats, regulatory compliance, and the unique challenges of healthcare data analysis. Users facing computational constraints are encouraged to explore cloud-friendly frameworks that leverage cloud computing services such as Azure and AWS. These platforms can alleviate the burden of substantial computational demands and the complexities of local infrastructure development. Moreover, cloud computing can significantly mitigate the risk of network issues that may arise from client-hosted infrastructures with varying capabilities. More specifically, we propose the following suggestions for FL framework selection, usage, and development:
  • Secure Communication: The integrity of the FL system hinges on secure communication protocols, where encryption should be employed 121.
  • Flexible Deployment: To streamline the deployment process, FL frameworks should support secure, reliable, and flexible deployment methods. They should integrate seamlessly with existing IT and data science infrastructures, facilitating a routine and uneventful deployment experience 122.
  • Data Science Workflow Support: Given the necessity for diverse data providers to achieve robust FL outcomes, frameworks should support a comprehensive data science workflow. The ideal framework should be agnostic to both the model and the data, accommodating a wide range of data types and analytical methods 80.
  • Scalability: Scalability is a key consideration for FL frameworks, which must accommodate an increasing number of participants and the corresponding complexity. Addressing scalability challenges, particularly with privacy-enhancing technologies such as synthetic data or HE, is crucial for the long-term viability of FL initiatives 123.
  • Future-proofing: FL frameworks should be designed with future-proofing in mind, anticipating emerging use cases, evolving security threats, and new privacy concerns. It should facilitate the dynamic participation of data providers, adapt computational resources to fluctuating client numbers, and implement regular system updates to address privacy and security challenges 122.

### v2

@@ 1 Introduction
The rest of this review is structured as follows. Section 2 provides an overview of the background and preliminaries of FL. Section 3 describes the screening procedure adopted in this work. Section 4 highlights the key findings of our analysis. Section 5 explores recent advances, challenges, and pitfalls in implementing FL in healthcare, offering practical recommendations to overcome current limitations and outlining potential future research directions.
@@ 2 Preliminaries
Beyond centralized or decentralized topologies, FL has evolved to address complex scenarios caused by varying feature and sample distributions. This evolution has led to the development of three primary paradigms: Horizontal Federated Learning (HFL), where data from different clients significantly overlap in the feature space but have little overlap in the sample space; Vertical Federated Learning (VFL), where data from different clients have minimal overlap in the feature space but significant overlap in the sample space; and Federated Transfer Learning (FTL), which leverages knowledge transfer to handle scenarios where there is little overlap in both feature and sample spaces. Figure 2 illustrates these differences. HFL is the most prevalent paradigm in FL studies. For instance, Clients 1 and 2 in Figure 2 represent scenarios where a vast number of people use wearables, such as the Apple Watch, to monitor their health conditions. These devices generate large amounts of data that share the same feature space (e.g., heart rate, step count), enabling collaborative model training. VFL, on the other hand, is better suited for applications where clients share the same sample space but store distinct features. For example, as illustrated by Clients 1 and 3 in Figure 2, a patient’s medical records may be distributed across multiple hospitals, with each hospital contributing unique features (e.g., imaging data, lab results). Aggregating these features allows for a more comprehensive model. Finally, FTL addresses scenarios with limited overlap in both feature and sample spaces. As depicted by Clients 2 and 4 in Figure 2, different clients may manage varying combinations of healthcare data and patient populations, with only a small intersection in the feature space. This approach is particularly relevant for tabular EHR data 28.
FL can be further categorized into Cross-silo FL and Cross-device FL based on the scale and attributes of participating clients. Cross-silo FL is tailored for scenarios where a limited number of participating clients, such as hospitals, medical centres, and institutions, collaboratively engage in all stages of FL training. Notable examples include Healthchain 12, which facilitates FL deployment among multiple hospitals in Europe, and the Melody Project 15, designed to optimize the drug discovery task across multiple companies while preserving data privacy. Cross-device FL, on the other hand, is designed for scenarios involving a multitude of participating clients, typically edge devices with limited data storage and computing capabilities. Examples include wearables (e.g., Apple Watch) and Internet of Medical Things (IoMT) devices. For instance, IoMT devices like Raspberry Pi and Jetson Nano can collect electronic health records (EHRs) in resource-limited environments, enabling early detection of sepsis 29.
@@ 3 Method
Our review was conducted in accordance with the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines 30. As shown in Figure 3, the flow diagram outlines the search, inclusion and exclusion procedures. We carried out a comprehensive search of the most recent studies focusing on advanced FL technologies within healthcare domain up to May 2024.
@@ 4.1 Application and Data
Included studies explored a broad range of healthcare specialties, including general medicine 6, cardiology 31, oncology 32, ophthalmology 33, drug discovery 15, multiomics 34, dermatology 35, and radiology 36. Most studies focused on tasks such as classification (67/107), segmentation (20/107), and detection (8/107), with additional applications in regression 37, clustering 34, reconstruction 38, 39, feature selection 40, 41, data synthesis 42, 43, 44, and biomedical language process 45. In terms of data types, medical imaging, including MRI, CT, and X-rays, was the most frequently used (55/107), followed by clinical data and electronic health records (EHR) (24/107), skin images (6/107), retinal images (6/107), histopathology slides (16/107), multiomics data (3/107), and biomedical language data (1/107). Some studies involved multiple data types, while 8 studies did not specify the type of data used or used non-healthcare data 46, 47, 48.
@@
Cohort Size $\leq 100$ $100-1000$ $\geq 1000$ unavailable 6 9 24 68 23, 49, 50 5, 6, 7 8, 37, 14 51, 52, 38
Task Classification Segmentation Detection Others 74 22 8 10 53, 54, 55 56, 57, 58 19, 29, 59 37, 34, 39
Data Type Medical Imaging (MRI, CT, X-rays) Clinical and EHR Skin Images Retinal Images Histopathology Slides Multiomics Biomedical Language unavailable or non-healthcare 55 24 6 6 16 3 1 8 60, 6, 8 61, 62, 63 51, 64, 65 66, 64, 58 24, 67, 68 34, 69, 70 45 71, 48, 47
Number of Clients $\leq 10$ $10-50$ $\geq 50$ unavailable 57 17 6 27 7, 11, 15 5, 8, 9 62, 71, 72 73, 74, 48
Topology Centralized Decentralized 95 12 53, 54, 38 75, 76, 46
Type of Federation Cross-Silo Cross-Device unavailable 98 7 2 77, 39, 58 76, 75, 78 48, 74
Scenario Deployment Simulation 10 97 79, 80, 81 82, 83, 65
Framework Custom-designed Open-source Options unavailable 78 13 16 83, 82, 84 13, 8, 80 52, 65, 85
Data Curation Standardization & Harmonization unavailable 12 95 7, 67, 23 51, 86, 87
Data Partition HFL VFL FTL unavailable 94 3 1 12 88, 89, 90 28, 50, 38 78 42, 91, 41
Natural Split Simulate unavailable 37 67 3 76, 92, 39 93, 45, 87 94, 80, 9
Train/Test/Val Details per Client 17 45, 90, 95
Holdout Cohort for Evaluation 12 51, 39, 67
Model Deep Learning Traditional Machine Learning unavailable 83 7 17 79, 81, 29 7, 37, 63 96, 13, 80
Initialization: Random Initialization: Pretrained/Foundation Models unavailable 18 5 84 7, 28, 66 97, 35, 45 98, 84, 55
@@
Generalization in Open Domain 15 7, 28, 35
Communication Efficiency 19 62, 27, 99
@@
Temporal Data Dynamics 2 29, 28
Synchronous Aggregation Asynchronous Aggregation 92 15 19, 23, 24 34, 35, 36
Privacy and Security Model Updates Protection 41 19, 100, 65
Open Source Code Available Trained Model Available unavailable 29 1 78 53, 54, 66 8 101, 49, 99
@@ 4.2 Topology, Scenario and Framework
The centralized FL paradigm dominates current implementations, with 95 out of 107 studies following this topology. Only 10 studies reported real-world deployments in distributed clinical settings, while the rest remained in the realm of prototypes or simulations. In terms of frameworks, the majority (78/107) utilized custom-designed FL frameworks, while a smaller number (13/107) employed open-source options such as Flower 13, Flare 8, 80, 11, SubstraFL 7, 15, TFF 37, OpenFL 14, PySyft 36, 49, 102, 67, and FedBioMed 81. 16 studies did not specify the framework used. Further details about open-source frameworks can be found in Table 2.
@@ 4.3 Data Curation and Partition
Among the reviewed studies, only 12 provided details on the processes of data standardization and harmonization. Regarding data partition, HFL was the predominant approach, with 94 out of 107 studies focusing solely on it. In contrast, VFL was explored in only 3 studies 38, 28, 50, while 2 studies considered both HFL and VFL in combination 28, 50. Only 1 study discussed FTL 78. Notably, 12 studies did not mention this aspect at all. The majority of studies addressed only one type of data heterogeneity, such as quantity skew or label skew, without considering multiple factors simultaneously. Moreover, 37 studies employed natural data splits for training and/or evaluation, while the rest relied on artificial splits. Only 17 studies detailed their training, testing or validation sets and 12 studies split a holdout cohort for evaluation.
@@ 4.4 Model
Among reviewed studies, Convolutional Neural Networks (CNNs) were the most commonly utilized (80/107), including both custom models specifically designed for healthcare tasks and well-established architectures like ResNet, DenseNet, MobileNet, and U-Net 101, 49, 99. Additionally, Recurrent Neural Networks (RNNs) have been incorporated to leverage their strengths in handling complex healthcare data 50, 29, 103. Some studies also employed custom Multi-Layer Perceptrons (MLPs) and attention mechanisms to further boost model performance 71, 104, 62. 45 utilized large language models for distributed biomedical natural language processing. Beyond deep learning approaches, several studies explored traditional machine learning (ML) algorithms, including linear models and ensemble methods. Notable examples include logistic regression 71, support vector machines 85, fuzzy clustering 105, and decision trees 72, 106. Interestingly, only 23 studies explicitly discussed their initialization strategies for model training. Among these, the majority opted for random initialization, while a mere five clearly stated that they utilized pretrained or foundation models as their starting point 97, 107, 35, 31, 45.
@@ 4.5 Optimization
Most studies addressed either data or model heterogeneity, and none of them considered system heterogeneity. Only 15 studies evaluated model generalization ability in unseen open domains. A total of 19 studies focused on improving communication efficiency, employing techniques such as knowledge distillation 53, 100, 97, gradient quantization 62, one-shot FL 99, split learning 102, and tensor factorization 96, 108. In terms of convergence analysis, a few studies (21/107) reported metrics such as communication rounds and costs, as well as overall convergence time, but none provided a theoretical understanding of convergence dynamics. Only two studies considered temporal data dynamics in model learning 28, 29. Regarding synchronization, 15 studies employed asynchronous aggregation instead of synchronous aggregation, particularly in applications involving wearables 76 and IoMT devices 29.
@@ 4.6 Privacy and Security
Only 41 studies addressed the exchange of model updates with privacy guarantees. The most commonly used techniques for safeguarding model updates included Differential Privacy (DP), Homomorphic Encryption (HE), Secure Multi-Party Computation (SMPC), knowledge distillation, and partial model exchange. However, metadata such as sample sizes and distributions were frequently shared without protection, particularly in methods based on FedAvg 107, 109. To mitigate the risk of adversaries inferring raw data, synthetic data was employed in some cases 19, 99, 44, 86. Additionally, swarm learning and blockchain were utilized to secure the communication process 24, 19.
@@ 4.7 Fairness and Incentive
Only three studies have discussed issues related to fairness and/or incentives in healthcare FL 110, 47, 55, with just one of these studies specifically exploring the complexities of both fairness and incentives in detail 55.
In the context of FL for healthcare, fairness generally refers to the equitable distribution of model performance among participants, ensuring that no entity is disadvantaged. Incentives are mechanisms designed to motivate healthcare institutions to participate in federated networks, often by offering rewards for contributions such as high-quality data or computational resources. For a more comprehensive discussion of fairness and incentive in healthcare FL, we refer readers to Section 5.6, where these concepts are explored in greater detail.
@@ 4.8 Evaluation
Most studies used conventional ML metrics for evaluation, such as accuracy, precision, area under the receiver operating characteristic curve (AUC), sensitivity/recall, specificity, F1-score, Dice score, Intersection over Union (IoU), Hausdorff Distance (HD), and loss value. Additionally, many studies performed comparisons against classical centralized models or localized models, and conducted ablation studies. However, only a few studies (26/107) addressed critical aspects unique to FL, such as communication overhead, resource consumption, scalability, generalization, privacy, fairness, and security concerns. As for benchmarking, just one study provided relatively comprehensive benchmarks across multiple healthcare datasets 54. Interpretability was explored in seven studies, either through feature selection 63, 40, attention maps 41, 53, 67, or tree-based models 7, 37. While 29 studies released their source code, only one also made the trained model publicly available 8.
@@ 5.1.1 Domination of Simulation Studies
Moreover, very few studies have explored FL practices at a national or international scale 8, 63. Notable examples include the Collaborative Data Analysis (CODA) 79, 15, HealthChain 12 and DRAGON 13. CODA tested FL’s feasibility across eight hospitals in Canada by enrolling patients with suspected or confirmed COVID-19 over three years. Melody deployed multi-task FL among 10 pharmaceutical companies to optimize the drug discovery process. HealthChain and DRAGON implemented FL across multiple hospitals in Europe, facilitating the prediction of treatment responses for cancer and COVID-19 patients.
@@
Flare 80 Nvidia DP, HE - -
FedML 111 TensorOpera DP, HE -
FederatedScope 112 Alibaba DP, SMPC - - - - - -
Flower 113 FlowerLab DP - -
FATE 114 WeBank DP, HE, SMPC - -
SubstraFL 115 Owkin DP, SMPC - -
PySyft 116 OpenMined DP - - - - - - -
OpenFL 117 Intel DP - - - - - -
TFF 118 Google DP - - - - - -
Fed-BioMed 81 Inria DP - - - - -
IBM FL 119 IBM DP, SMPC - - - - - -
PaddleFL 120 Baidu DP, SMPC - - - - - -
SAFEFL 121 ENCRYPTO SMPC - - - - - - -
@@ 5.1.2 Deficiency in FL Frameworks Development and Usage
  • For researchers aiming to swiftly prototype and test novel concepts can benefit from frameworks that incorporate comprehensive simulator modules. These tools allow for the rapid iteration and validation of ideas within a controlled simulation environment, which can be critical for the initial stages of research and development.
  • For engineers seeking to deploy FL in real-world scenarios should consider frameworks tailored to the specific needs of the healthcare domain. These frameworks should offer healthcare specific adaptations to ensure compatibility with medical data formats, regulatory compliance, and the unique challenges of healthcare data analysis.
  • Users facing computational constraints are encouraged to explore cloud-friendly frameworks that leverage cloud computing services such as Azure and AWS. These platforms can alleviate the burden of substantial computational demands and the complexities of local infrastructure development. Moreover, cloud computing can significantly mitigate the risk of network issues that may arise from client-hosted infrastructures with varying capabilities.
  • More specifically, we propose the following suggestions for FL framework selection, usage, and development:
    • Secure Communication: The integrity of the FL system hinges on secure communication protocols, where encryption should be employed 122.
    • Flexible Deployment: To streamline the deployment process, FL frameworks should support secure, reliable, and flexible deployment methods. They should integrate seamlessly with existing IT and data science infrastructures, facilitating a routine and uneventful deployment experience 123.
    • Data Science Workflow Support: Given the necessity for diverse data providers to achieve robust FL outcomes, frameworks should support a comprehensive data science workflow. The ideal framework should be agnostic to both the model and the data, accommodating a wide range of data types and analytical methods 81.
    • Scalability: Scalability is a key consideration for FL frameworks, which must accommodate an increasing number of participants and the corresponding complexity. Addressing scalability challenges, particularly with privacy-enhancing technologies such as synthetic data or HE, is crucial for the long-term viability of FL initiatives 124.
    • Future-proofing: FL frameworks should be designed with future-proofing in mind, anticipating emerging use cases, evolving security threats, and new privacy concerns. It should facilitate the dynamic participation of data providers, adapt computational resources to fluctuating client numbers, and implement regular system updates to address privacy and security challenges 123.